OpenStudy (anonymous):

Write product in simplest form: (1 - 1/2) (1 - 1/3) (1 - 1/4) ... (1-1/1000)

OpenStudy (anonymous):

(1 - 1/2) (1 - 1/3) (1 - 1/4) ... (1-1/1000) = (1/2)(2/3)(3/4)(4/5) ... (998/999)(999/1000) = 1/1000